Two cards are drawn one after the other with replacement from a well-shuffled pack of card. What is the probability that the two cards are both kings?
Found 2 solutions by Edwin McCravy, mccravyedwin:
Answer by Edwin McCravy(20077)


(4/52)(4/52) = (1/13)(1/13) = 1/169

Edwin

Answer by mccravyedwin(421)

Probability is the fraction of times you would succeed at the activity if you
repeated the activity many times.

4/52 = 1/13 = 7.7% of the times you drew the first card you would get a king. 
But only 7.7% of those times when you drew a king the first time would you also
get a king on the second draw.

Let's suppose you drew two cards with replacement every day for 1000 days. Then
on 7.7% of those days, or about 77 days you would draw a king the first time.
But of those 77 days when you drew a king the first time, only 7.7% of those 77
days, or about 6 days would you draw two kings.

Answer = (4/52)(4/52) = (1/13)(1/13) = 1/169.  Only about 1 time out of 169
tries would you get two kings.  Maybe not exactly 1 time out of every 169 tries,
but on the average you would. 

Edwin